OpenStudy (anonymous):

If Sandy can afford car payments of $260 per month for 7 years, what is the price of a car she can afford now? Assume an interest rate of 6.6%. Any help is appreciated :)

jimthompson5910 (jim_thompson5910):

Total Amount = (Monthly Payment)*(Number of months in a year)*(# of years) Total Amount = (260)*(12)*(7) Total Amount = 21,840 So the amount she can afford is 21,840. This is the amount of the original value of the car + interest Assuming we're talking about simple interest, then we use the formula. A = P(1 + rt) So plug in A = 21840, r = 0.066, and t = 7 and solve for P 21840 = P(1 + 0.066*7) 21840 = P(1.462) 21840/1.462 = P P = 14,938.440492476 P = 14,938.44 So she can afford a $14,938.44 car.

OpenStudy (anonymous):

Thanks you so much :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

I assume you know the Present value formula to apply for annuity payment type!